A box of mass 4.5kg is pull at a constant speed of 2m/s along a rough horizontal floor by a horizontal force of magnitude 15N.
(i) find the coefficient of friction between the box and the floor. The horizontal pulling force is now removed. Find:
(ii) the acceleration of the box in the subsequent motion.
1
Expert's answer
2019-10-24T09:34:27-0400

(i) Since the speed is constant, sum of the force of the friction and pulling force is 0.

So, "15=k*4.5*9.8" .

Then, "k=\frac{15}{4.5*9.8}=0.34."


(ii) "f=ma,\;\;a=\frac{f}{m}=\frac{-15}{4.5}=-3.33\; m*s^{-2}"